The hourly wage y of an employee at a certain production company is given by y=0.25x + 9 where x is the number of units produced in an hour.
a) complete the table and graph the results
X 0 1 5 10
Y
B) Find the number of units thatmust be produced each hour to earn an hourly wage of $12.25.

Answer by elima(1433)

The numbers given to you for, you would plug those into the equation for x;
y=0.25x+9
y=0.25(0)+9
y=9
So your first point would be; (0,9)
y=0.25(1)+9
y=0.25+9
y=9.25
Point 2; (1,9.25)
y=0.25(5)+9
y=1.25+9
y=10.25
point 3; (5,10.25)
y=0.25(10)+9
y=2.5+9
y=11.25
point 4; (10,11.25)
So the number of units produced to earn $12.25 per hour is going to be more than 10;
y=0.25(13)+9
y=3.25+9
y=12.25
So the number of units would be 13
=)